Question 1: Cost Minimization and Cost Curves Suppose that Jennifer produces goody by using input x and x2. The production fuMicroeconomics 2 (b) Now that weve worked out the necessary condition for cost minimisation, use this condition to find the(c) Now that we have found the conditional demand, substitute them back into the cost function. Find the Average Cost, Averag=For parts (d) and (e) assume F=6, wi=2 and wz=4. (d) i) Sketch the AC, AVC, AFC and MC curves on the grid below which uses the) Show that the MC curve intersects the AVC curve and the AC curve at the minimum point.
Answer #1

me Given that in Production function f(x,, &2) =Y= 5x + 2 cost funchon c=wili+W, 2q + F cost minimisa kon problemi- minmisatiPage 2 ered by ime ho Necesory contion conditional (B? X, (co,W2 ,y) and & Co. ,W2, Y) are the input demand functions 9: 7, +Page? 3 . Wiekted X2 dx2 door X2 = y?w7 - , t W2 + 2W, W2 dX2 , 2017 W24200,00) Cay?co,)-(1+2w2]y?? (co, + O2 + 20,W2)² diz 2Pogethe - Tc. = 2y²Cu32 + uy? ( 28²th retruje cret Full » TC= 32 tilan to > TC - - Log? 8134276 to = 8/34²76 → AC = 8/3y76)